# Convergence Rates of Constrained Expected Improvement

### v1

@@ 4 Numerical examples
Next, we evaluate simple regret of CEI on five benchmark examples. All these benchmark examples are chosen from the most commonly used test problems in CBO. Specifically, Example 3 tests performance in a small feasible region, which was previously studied in [1, 30]. Example 4 includes multiple constraints and local minimums, which has been used in [31, 32, 33, 30]. Example 5 is a four-dimensional problem, previously studied in [32, 30]. Example 6 is the six-dimensional Hartmann problem, previously tested in [34]. Example 7 is the so-called banana function, where the global minimum lies in a narrow region. The mathematical formulations of Examples 3–7 are presented in Section D in the appendix. For two-dimensional problems, we also include the contour plots of the objective and constraint functions in Appendix D.
Each test problem is repeated for 100 independent experiments. The hyper-parameters of the kernels are estimated by a standard maximum likelihood method. When CEI fails to identify a feasible sample, we adopt the same heuristic strategy as in [34]. The numerical results are summarized in Figure 1(c) - Figure1(g). The solid line represents the median of the simple regret, and the dotted lines represent the 25th percentile and 75th percentile of the simple regret, respectively.
@@
From the figures, we observe that CEI consistently reduces simple regret, aligning with the asymptotic convergence theories established in this paper. For each of the seven examples, the simple regret converges to 00 quickly. The 25th percentile and 75th percentile results demonstrate the good statistical properties of CEI.
@@ 5 Conclusions
In this paper, we studied the simple regret upper bounds of the CEI algorithm, one of the most widely adopted CBO methods. Under both frequentist and Bayesian assumptions on the objective and constraint functions, we establish for the first time the convergence rates for CEI. Our results provide theoretical support and validation for the empirical success of CEI.
@@ A.1 Information gain
To obtain state-of-the-art simple regret upper bound, we utilize the information theory results that are well-established in previous literature [28, 29, 25]. Using $f$ as the example, let $A\subset C$ denote a set of sampling points. Assume that the observations are noisy at these points are noisy with $y_{A}=f({\bm{x}}_{A})+\epsilon_{A}$ at ${\bm{x}}\in A$, where $\epsilon_{A}\sim\mathcal{N}(0,\sigma^{2})$ denotes the independent and identically distributed Gaussian noises. The maximum information gain is defined as follows.
@@
The rate of increase for $\gamma^{f}_{t}$ depends on the property of the kernel. For common kernels such as the SE kernel and the Matérn kernel, $\gamma^{f}_{t}$ has been widely studied in literature and the state-of-the-art rates of $\gamma^{f}_{t}$ are summarized in Lemma A.2 [25].
@@
For GP with $t$ samples, the SE kernel has $\gamma_{t}=\mathcal{O}(\log^{d+1}(t))$, and the Matérn kernel with smoothness parameter $\nu>0$ has $\gamma_{t}=\mathcal{O}(t^{\frac{d}{2\nu+d}}(\log^{\frac{2\nu}{2\nu+d}}(t)))$.
@@
While $\gamma_{t}^{f}$ is defined in the noisy case, we can readily apply it to the noise-free case and bound $\sigma_{t}^{f}({\bm{x}}_{t+1})$ using techniques similar to that in [22]. To do so, we note that given the Gaussian observation noise $\epsilon_{t}\sim\mathcal{N}(0,\sigma^{2})$, we can write the posterior prediction for $f$ as
@@
Similarly, we can define the posterior predictions for $c$ as $\tilde{\mu}_{t}^{f}({\bm{x}})$ and $\tilde{\sigma}_{t}^{f}({\bm{x}})$. The sum of posterior variance for GP generated by (2) satisfy the next lemma, based on information theory [28] .
@@
The noise-free posterior standard deviation satisfies $\sigma_{t}^{f}({\bm{x}})<\tilde{\sigma}_{t}^{f}({\bm{x}})$ for $\forall\sigma>0$.
@@
The posterior standard deviation under assumptions (1)-(4) in [16] in the frequentist and noise-free setting is given in Lemma 3.6, whose proof is given below.
@@
Both the SE and Matérn kernels satisfy assumptions (1)-(4) in [16]. By Lemma 7 in [16], there exists $C^{\prime}>0$ so that
@@
holds for at most $k$ times, for $\forall k\leq t$. Similarly, for Matérn kernels,
@@
at most $k$ times. For Matérn kernels,
@@
The Lipschitz continuity of function $f$ extends to the maximum function $f^{+}=\max\{f,0\}$ in the following lemma.
Lemma A.6.
If a function $f:\mathbb{R}^{n}\to\mathbb{R}$ is Lipschitz continuous on $O\subset\mathbb{R}^{n}$ with Lipschitz constant $L$ in 1-norm, then its maximum with 00 is also Lipschitz, i.e.
$|f^{+}({\bm{x}})-f^{+}({\bm{x}}^{\prime})|\leq L\left\lVert{{\bm{x}}-{\bm{x}}^{\prime}}\right\rVert_{1},$ (30)
where $f^{+}({\bm{x}})=\max\{f({\bm{x}}),0\}$ and ${\bm{x}},{\bm{x}}^{\prime}\in O$.
Proof.
From the definition of Lipschitz functions, we know
$|f({\bm{x}})-f({\bm{x}}^{\prime})|\leq L\left\lVert{{\bm{x}}-{\bm{x}}^{\prime}}\right\rVert_{1},$ (31)
If $f({\bm{x}})>0$ and $f({\bm{x}}^{\prime})>0$, then
$|f^{+}({\bm{x}})-f^{+}({\bm{x}}^{\prime})|=|f({\bm{x}})-f({\bm{x}}^{\prime})|\leq L\left\lVert{{\bm{x}}-{\bm{x}}^{\prime}}\right\rVert_{1}.$ (32)
If $f({\bm{x}})>0$ and $f({\bm{x}}^{\prime})\leq 0$, then
$|f^{+}({\bm{x}})-f^{+}({\bm{x}}^{\prime})|=f({\bm{x}})\leq f({\bm{x}})-f({\bm{x}}^{\prime})\leq|f({\bm{x}})-f({\bm{x}}^{\prime})|\leq L\left\lVert{{\bm{x}}-{\bm{x}}^{\prime}}\right\rVert_{1}.$ (33)
Similarly, if $f({\bm{x}})\leq 0$ and $f({\bm{x}}^{\prime})>0$, then
$|f^{+}({\bm{x}})-f^{+}({\bm{x}}^{\prime})|=f({\bm{x}}^{\prime})\leq f({\bm{x}}^{\prime})-f({\bm{x}})\leq|f({\bm{x}})-f({\bm{x}}^{\prime})|\leq L\left\lVert{{\bm{x}}-{\bm{x}}^{\prime}}\right\rVert_{1}.$ (34)
Finally if $f({\bm{x}})\leq 0$ and $f({\bm{x}}^{\prime})\leq 0$, then
$|f^{+}({\bm{x}})-f^{+}({\bm{x}}^{\prime})|=0\leq|f({\bm{x}})-f({\bm{x}}^{\prime})|\leq L\left\lVert{{\bm{x}}-{\bm{x}}^{\prime}}\right\rVert_{1}.$ (35)
@@
The last statement in Lemma A.7 is a well-known result (e.g., see proof of Lemma 5.1 in [28]).
@@
For $\forall{\bm{x}}\in C$, $EI_{t}^{f}({\bm{x}})\geq 0$ and $EI_{t}^{f}({\bm{x}})\geq y^{+}_{t}-\mu_{t}^{f}({\bm{x}})$. Moreover,
$z_{t}^{f}({\bm{x}})\leq\frac{EI_{t}^{f}({\bm{x}})}{\sigma_{t}^{f}({\bm{x}})}<\begin{cases}\phi(z_{t}^{f}({\bm{x}})),\ &z_{t}^{f}({\bm{x}})<0,\\ z_{t}^{f}({\bm{x}})+\phi(z_{t}^{f}({\bm{x}})),\ &z_{t}^{f}({\bm{x}})\geq 0.\end{cases}$ (36)
@@ Appendix B Proofs for simple regret upper bound under frequentist assumptions
The following Lemma is a well-established result [29].
@@
At any given ${\bm{x}}\in C$ and $t\in\mathbb{N}$, the prediction errors satisfy
@@
Next, we present a lemma on the relationship between $I_{t}^{f}$ and $EI_{t}^{f}$, previously seen in [16].
@@
$P_{t_{k}}({\bm{x}}^{*})=\Phi\left(\frac{-\mu_{t_{k}}^{c}({\bm{x}}^{*})}{\sigma_{t_{k}}^{c}({\bm{x}}^{*})}\right)\geq\Phi(-B_{c}).$ (55)
@@
$r_{t}\leq$ $\frac{c_{\tau B}}{\Phi(-B_{c})}\left[\frac{2M_{f}}{k}+(B_{f}+0.4)\sigma_{t_{k}}^{f}({\bm{x}}_{t_{k}+1})\right].$ (56)
As $t\to\infty$, $t_{k}\to\infty$ and $k\to\infty$. Further, if $\sigma_{t}({\bm{x}}_{t+1})\to 0$, $\sigma_{t_{k}}^{f}({\bm{x}}_{t_{k}+1})\to 0$. ∎
Proof of Theorem 3.7 is presented next.
@@
$r_{t}\leq$ $\frac{c_{B_{f}}}{\Phi(-B_{c})}\left[\frac{2M_{f}}{k}+(B_{f}+0.4)C^{\prime}k^{-\frac{\min\{\nu,1\}}{d}}\log^{\eta}(k)\right].$ (57)
@@ B.1 Proofs for improved simple regret upper bound under frequentist assumptions
Appendix C Proofs for simple upper bound regret under Bayesian assumptions
@@
Under Assumption 3.10 and 3.11, there exists $M_{f}>0$ such that $|f({\bm{x}})|\leq M_{f}$. Similarly, there exists $M_{c}>0$ such that $|c({\bm{x}})|\leq M_{c}$.
We recall a well-known result on the bounds of $|f({\bm{x}})-\mu_{t}^{f}({\bm{x}})|$ under Assumption 3.10 [28].
@@
holds with probability $\geq 1-\delta$. Similarly, with $\delta\in(0,1)$ and $\beta=2\log(1/\delta)$.
$|c({\bm{x}})-\mu_{t}^{c}({\bm{x}})|\leq\sqrt{\beta}\sigma_{t}^{c}({\bm{x}}),$ (60)
with probability $\geq 1-\delta$.
@@
Let $e^{-\frac{\beta}{2}}=\delta$ and (59) is proven. Similarly, (60) can be proven. ∎
@@
Under Assumption 3.10, the probability distribution of $I_{t}$ satisfies
$\mathbb{P}\{I_{t}^{f}({\bm{x}})\leq a\}=\begin{cases}0,\ &a<0,\\ \Phi\left(\frac{a}{\sigma_{t}^{f}({\bm{x}})}-z_{t}^{f}({\bm{x}})\right),\ &a\geq 0.\end{cases}$ (64)
@@
Next, we present the relationship between $I_{t}^{f}({\bm{x}})$ and $EI_{t}^{f}({\bm{x}})$. Lemma C.4 was previously first presented and proved in [35].
@@
For the second probability in (66), we further distinguish between two cases. First, consider $-\sigma_{t}^{f}({\bm{x}})w+EI_{t}^{f}({\bm{x}})<0$. From Lemma C.3,
@@
Second, consider the premise $-\sigma_{t}^{f}({\bm{x}})w+EI_{t}^{f}({\bm{x}})\geq 0$. By Lemma C.3, we have
@@
with probability greater than $21-\delta/$ 2. Using $\delta/2$ in Lemma C.4,
@@
$\sum_{i=0}^{t-1}f^{+}_{i}-f^{+}_{i+1}=f^{+}_{0}-f_{t}^{+}\leq 2M_{f}.$ (87)
Given that $f^{+}_{i}-f^{+}_{i+1}\geq 0$, $f^{+}_{i}-f^{+}_{i+1}\geq\frac{2M_{f}}{k}$ at most $k$ times for any $k\in\mathbb{N}$. Choose $k=[t/2]$, where $[x]$ is the largest integer smaller than $x$ so that $2k\leq t\leq 2(k+1)$. Then, there exists $k\leq t_{k}\leq 2k$ so that $f_{t_{k}}^{+}-f_{t_{k}+1}^{+}<\frac{2M_{f}}{k}$ and $f({\bm{x}}_{t_{k}+1})\geq f_{t_{k}+1}^{+}$.
@@
where the last inequality uses $\phi(\cdot)<0.4$. From Lemma C.2, we have with probability $\geq 1-\delta/3$,
$f^{+}_{t_{k}}-\mu_{t_{k}}^{f}({\bm{x}}_{t_{k}+1})=$ $f^{+}_{t_{k}}-f({\bm{x}}_{t_{k}+1})+f({\bm{x}}_{t_{k}+1})-\mu_{t_{k}}^{f}({\bm{x}}_{t_{k}+1})$ (89)
$\leq$ $f^{+}_{t_{k}}-f^{+}_{t_{k}+1}+\beta^{1/2}\sigma_{t_{k}}^{f}({\bm{x}}_{t_{k}+1})\leq\frac{2M_{f}}{k}+\beta^{1/2}\sigma_{t_{k}}^{f}({\bm{x}}_{t_{k}+1}).$
Applying (89) to (88), we have with probability $\geq 1-2\delta/3$,
$r_{t}\leq$ $c_{\tau}(\beta)\frac{P_{t_{k}}({\bm{x}}_{t_{k}+1})}{P_{t_{k}}({\bm{x}}^{*})}\left[\frac{2M_{f}}{k}+(0.4+\beta^{1/2})\sigma_{t_{k}}^{f}({\bm{x}}_{t_{k}+1})\right].$ (90)
Next, we consider the probability function $P_{t_{k}}$ at ${\bm{x}}^{*}$ and ${\bm{x}}_{t_{k}+1}$. Using the fact that $c({\bm{x}}^{*})\leq 0$, we have by Lemma C.2 at ${\bm{x}}^{*}$ and $t_{k}$,
$\mu_{t_{k}}^{c}({\bm{x}}^{*})\leq\sqrt{\beta}\sigma_{t_{k}}^{c}({\bm{x}}^{*})+c({\bm{x}}^{*})\leq\sqrt{\beta}\sigma_{t_{k}}^{c}({\bm{x}}^{*}),$ (91)
with probability $\geq 1-\delta/3$. Thus, with probability $\geq 1-\delta/3$,
$\frac{-\mu_{t_{k}}^{c}({\bm{x}}^{*})}{\sigma_{t_{k}}^{c}({\bm{x}}^{*})}\geq-\sqrt{\beta}.$ (92)
From the monotonicity of $\Phi$, we have with probability $\geq 1-\delta/3$,
$\Phi\left(\frac{-\mu_{t_{k}}^{c}({\bm{x}}^{*})}{\sigma_{t_{k}}^{c}({\bm{x}}^{*})}\right)\geq\Phi\left(-\sqrt{\beta}\right),$ (93)
@@
$\frac{P_{t_{k}}({\bm{x}}_{t_{k}+1})}{P_{t_{k}}({\bm{x}}^{*},\epsilon)}\leq\frac{1}{\Phi(-\sqrt{\beta})},$ (94)
with probability $\geq 1-\delta/3$. Applying (94) to (90), we have
$r_{t}\leq$ $c_{\tau}(\beta)\frac{1}{\Phi(-\sqrt{\beta})}\left[\frac{2M_{f}}{k}+(0.4+\beta^{1/2})\sigma_{t_{k}}^{f}({\bm{x}}_{t_{k}+1})\right],$ (95)
with probability $\geq 1-\delta$. As $t\to\infty$, $t_{k}\to\infty$ and $k\to\infty$.
@@
From Lemma A.5, $\sigma_{i}^{f}({\bm{x}}_{i+1})\geq\frac{\sqrt{\gamma_{t}^{f}t}}{k}$, where $i=1,\dots,t-1$, at most $k$ times for any $k\in\mathbb{N}$ and $k\leq t$. Choose $k=[t/3]$ which leads to $3k\leq t\leq 3(k+1)$. There exists $t_{k}$ so that $\sigma_{t_{k}}^{f}({\bm{x}}_{t_{k}+1})\leq\frac{\sqrt{t\gamma^{f}_{t}}}{k}$. Following the proof for (95), we have
$r_{t}\leq$ $c_{\tau}(\beta)\frac{1}{\Phi(-\sqrt{\beta})}\left[\frac{2M_{f}}{k}+(0.4+\beta^{1/2})\frac{\sqrt{t\gamma^{f}_{t}}}{k}\right],$ (96)
@@
Appendix D Example details
The mathematical formulations of the testing problems are given below. The objective, constraint functions and the optimal $f$ of example 3 is given below.

### v2

@@ 4.2 Test problems
Next, we evaluate simple regret of CEI on five commonly used test problems in the literature of CBO. Specifically, Problem 1 tests performance in a small feasible region, which was previously studied in Gardner et al. (2014); Ariafar et al. (2019). Problem 2 includes multiple constraints and local minimums, which has been used in Gramacy et al. (2016); Hernández-Lobato et al. (2015). Problem 3 is a four-dimensional problem, previously studied in Picheny et al. (2016); Ariafar et al. (2019). Problem 4 is the six-dimensional Hartmann problem, previously tested in Letham et al. (2019). Problem 5 is the Rosenbrock function, where the global minimum lies in a narrow region. The mathematical formulations of the five functions are presented in Appendix D. For two-dimensional problems, we also include the contour plots of the objective and constraint functions in Appendix D. The SE kernel is used for the GP modeling (similar performance is observed for the Matérn kernel) and the hyper-parameters are estimated by a standard maximum likelihood method.
@@
For each test problem, we conducted 100 independent trials with different random initial designs. When CEI fails to identify a feasible sample, we adopt the same heuristic strategy as in Letham et al. (2019). The numerical results are summarized in Figure 2. The solid line represents the median of the simple regret, and the dotted lines represent the 25th percentile and 75th percentile of the simple regret, respectively. From the figures, we observe that CEI consistently reduces simple regret, aligning with the asymptotic convergence theories established in this paper. Accross all problems, the simple regret converges to 0 quickly. The 25th percentile and 75th percentile results demonstrate the good statistical properties of CEI.
(a) Problem 1: a small feasible region problem
(b) Problem 2: a multiple-constraints problem
(c) Problem 3: a four-dimensional problem
(d) Problem 4: Hartmann6 function
(e) Problem 5: Rosenbrock function (in log scale due to large range of objective values.)
Figure 2: Simple regret of CEI for five test problems.
@@ 5 Conclusions
In this paper, we studied the simple regret upper bounds of the CEI algorithm, one of the most widely adopted CBO methods. Under both frequentist setting and Bayesian objective assumptions, we establish for the first time the convergence rates for CEI. Our results provide theoretical support and validation for the empirical success of CEI.
@@ A.1 Information gain
To obtain state-of-the-art simple regret upper bound, we utilize the information theory results that are well-established in previous literature [Srinivas et al., 2009, Chowdhury and Gopalan, 2017, Vakili et al., 2021]. Using $f$ as the example, let $A\subset C$ denote a set of sampling points. Assume that the observations are noisy at sample points with $y_{A}=f({\bm{x}}_{A})+\epsilon_{A}$ at ${\bm{x}}\in A$, where $\epsilon_{A}\sim\mathcal{N}(0,\sigma^{2})$ denotes the independent and identically distributed Gaussian noises. The maximum information gain is defined as follows.
@@
The rate of increase for $\gamma^{f}_{t}$ depends on the property of the kernel. For common kernels such as the SE kernel and the Matérn kernel, $\gamma^{f}_{t}$ has been studied in literature and the state-of-the-art rates of $\gamma^{f}_{t}$ are summarized in Lemma A.2 [Vakili et al., 2021, Iwazaki, 2025b].
@@
For GP with $t$ samples, the SE kernel has $\gamma_{t}=\mathcal{O}(\log^{d+1}(t))$, and the Matérn kernel with smoothness parameter $\nu>0.5$ has $\gamma_{t}=\mathcal{O}(t^{\frac{d}{2\nu+d}}(\log^{\frac{2\nu}{2\nu+d}}(t)))$.
@@
While $\gamma_{t}^{f}$ is defined in the noisy case, we can readily apply it to the noise-free case and bound $\sigma_{t}^{f}({\bm{x}}_{t+1})$ using techniques similar to that in Lyu et al. [2019]. To do so, we note that given the Gaussian observation noise $\epsilon_{t}\sim\mathcal{N}(0,\sigma^{2})$ in the GP model, the posterior prediction for $f$ becomes
@@
Similarly, we can define the posterior predictions for $c$ with noise in the GP model as $\tilde{\mu}_{t}^{f}({\bm{x}})$ and $\tilde{\sigma}_{t}^{f}({\bm{x}})$. The sum of posterior variance for GP generated by (2) satisfy the next lemma, based on information theory Srinivas et al. [2009] .
@@
The noise-free (GP) posterior standard deviation satisfies $\sigma_{t}^{f}({\bm{x}})<\tilde{\sigma}_{t}^{f}({\bm{x}})$ for $\forall\sigma>0$.
@@
The posterior standard deviation under assumptions (1)-(4) in Bull [2011] in the frequentist and noise-free setting is given in Lemma 3.5, whose proof is given below.
@@
By Lemma 7 in Bull [2011], there exists $C^{\prime}>0$ so that
@@
holds for at most $k$ times, for $\forall k\leq t$. Similarly, for Matérn kernels with $\nu>0.5$,
@@
at most $k$ times. For Matérn kernels with $\nu>0.5$,
@@
The last statement in Lemma A.6 is a well-known result (e.g., see proof of Lemma 5.1 in Srinivas et al. [2009]).
@@
For $\forall{\bm{x}}\in C$, $EI_{t}^{f}({\bm{x}})\geq 0$ and $EI_{t}^{f}({\bm{x}})\geq f^{+}_{t}-\mu_{t}^{f}({\bm{x}})$. Moreover,
[…]
@@ Appendix B Proofs for simple regret upper bound under frequentist assumptions
The following Lemma is a well-established result Chowdhury and Gopalan [2017].
@@
At any given ${\bm{x}}\in C$ and $t\in\mathbb{N}$, the confidence intervals satisfy
@@
Next, we present a lemma on the relationship between $I_{t}^{f}$ and $EI_{t}^{f}$, previously seen in Bull [2011].
@@
$P_{t_{k}}({\bm{x}}^{*})=\Phi\left(\frac{-\mu_{t_{k}}^{c}({\bm{x}}^{*})}{\sigma_{t_{k}}^{c}({\bm{x}}^{*})}\right)\geq\Phi\left(-B_{c}\right).$ (48)
@@
$r_{t}\leq$ $\frac{c_{\tau B}}{\Phi(-B_{c})}\left[\frac{2B_{f}}{k}+(B_{f}+0.4)\sigma_{t_{k}}^{f}({\bm{x}}_{t_{k}+1})\right].$ (49)
As $t\to\infty$, $t_{k}\to\infty$ and $k\to\infty$. Further, if $\sigma_{t}^{f}({\bm{x}}_{t+1})\to 0$, $\sigma_{t_{k}}^{f}({\bm{x}}_{t_{k}+1})\to 0$ and $r_{t}\to 0$. ∎
Proof of Corollary 3.6 is presented next.
@@
$r_{t}\leq$ $\frac{c_{B_{f}}}{\Phi(-B_{c})}\left[\frac{2B_{f}}{k}+(B_{f}+0.4)C^{\prime}k^{-\frac{\min\{\nu,1\}}{d}}\log^{\eta}(k)\right].$ (50)
@@ B.1 Proofs for improved simple regret upper bound under frequentist assumptions
We provide the sample complexity of Theorem 3.7 below.
Corollary B.5.
Under Assumption 3.2, the CEI algorithm achieves a $\epsilon$ sample complexity of
$\mathcal{O}\!\left(\frac{1}{\epsilon^{2}}[\log(1/\epsilon)]^{d+1}\right)\ \text{and }\ \mathcal{O}\!\left(\epsilon^{-\frac{2\nu+d}{\nu}}\,\log(1/\epsilon)\right),$ (52)
for SE and Matérn kernels, respectively.
Proof.
Using Theorem 3.7, to achieve simple regret of most $\epsilon$ for SE kernel, set
$t^{-\frac{1}{2}}\log^{\frac{d+1}{2}}(t)=\epsilon.$
Solving asymptotically for $t$ gives the sample complexity
$t(\epsilon)=\mathcal{O}\!\left(\frac{1}{\epsilon^{2}}[\log(1/\epsilon)]^{d+1}\right).$
Similarly, for Matérn kernels, we have
$\epsilon=t^{\frac{-\nu}{2\nu+d}}\log^{\frac{\nu}{2\nu+d}}(t),$ (53)
which completes the proof. ∎
Appendix C Proofs for simple regret upper bound under Bayesian objective assumption
@@
Under Assumption 3.9, there exists $M_{f}>0$ such that $|f({\bm{x}})|\leq M_{f}$ with probability $\geq 1-\delta/3$. The constraint function is bounded by its RKHS norm bound $|c({\bm{x}})|\leq B_{c}$.
We recall a well-known result on confidence interval of $|f({\bm{x}})-\mu_{t}^{f}({\bm{x}})|$ under Assumption 3.9 Srinivas et al. [2009].
@@
holds with probability $\geq 1-\delta$.
@@
Let $e^{-\frac{\beta}{2}}=\delta$ and (54) is proven. ∎
Lemma C.3.
Given $\delta\in(0,1)$, let $\beta=2\log(\pi_{t}/\delta)$, where $\pi_{t}=\frac{\pi^{2}t^{2}}{6}$. Then, for all $t\in\mathbb{N}$,
$|f({\bm{x}})-\mu_{t}^{f}({\bm{x}})|\leq\sqrt{\beta_{t}}\sigma_{t}^{f}({\bm{x}}),$ (58)
holds with probability $\geq 1-\delta$.
@@
Under Assumption 3.9, the probability distribution of $I_{t}^{f}$ satisfies
[…]
@@
Next, we present the relationship between $I_{t}^{f}({\bm{x}})$ and $EI_{t}^{f}({\bm{x}})$.
@@
For the second probability in (61), we further distinguish between two cases. First, consider $\sigma_{t}^{f}({\bm{x}})w+EI_{t}^{f}({\bm{x}})<0$. From Lemma C.4,
@@
Second, consider the premise $\sigma_{t}^{f}({\bm{x}})w+EI_{t}^{f}({\bm{x}})\geq 0$. By Lemma C.4, we have
@@
with probability greater than $1-\delta/2$. Using $\delta/2$ in Lemma C.5,
@@
$\sum_{i=0}^{t-1}f^{+}_{i}-f^{+}_{i+1}=f^{+}_{0}-f_{t}^{+}\leq 2M_{f},$ (82)
with probability $\geq 1-\delta/3$. Next, consider $f_{i}^{+}-\mu_{i}^{f}(x_{i+1})$. Recall that $\beta_{t}=2\log(3\pi_{t}/\delta)$. From Lemma C.3 Lemma A.4 and Lemma A.3, we have
$\sum_{i=0}^{t-1}\max\{f_{i}^{+}-\mu_{i}^{f}({\bm{x}}_{i+1}),0\}=$ $\sum_{i=0}^{t-1}\max\{f^{+}_{i}-f({\bm{x}}_{i+1})+f({\bm{x}}_{i+1})-\mu_{i}^{f}({\bm{x}}_{i+1}),0\}$ (83)
$\leq$ $\sum_{i=0}^{t-1}f^{+}_{i}-f^{+}_{i+1}+\beta^{1/2}_{t}\sigma_{i}^{f}({\bm{x}}_{i+1})\leq 2M_{f}+\beta^{1/2}_{t}\sqrt{C_{\gamma}t\gamma_{t}^{f}},$
with probability $\geq 1-2\delta/3$ via union bound. Given that $\max\{f_{i}^{+}-\mu_{i}^{f}({\bm{x}}_{i+1}),0\}\geq 0$, $\max\{f_{i}^{+}-\mu_{i}^{f}({\bm{x}}_{i+1}),0\}\geq\frac{2M_{f}}{k}+\frac{\beta^{1/2}_{t}}{k}\sqrt{C_{\gamma}t\gamma_{t}^{f}}$ at most $k$ times for any $k\in\mathbb{N}$ with probability $\geq 1-2\delta/3$. Choose $k=[t/2]$, where $[x]$ is the largest integer smaller than $x$ so that $2k\leq t\leq 2(k+1)$. Then, choose the first index $t_{k}$ where $k\leq t_{k}\leq 2k$ so that $\max\{f_{t_{k}}^{+}-\mu_{t_{k}}^{f}({\bm{x}}_{t_{k}+1}),0\}<\frac{2M_{f}}{k}+\frac{\beta^{1/2}_{t}}{k}\sqrt{C_{\gamma}t\gamma_{t}^{f}}$. As discussed above, such a $t_{k}$ exists with probability $\geq 1-2\delta/3$. We note that maximum information gain and its upper bound does not depend on the optimization path. Importantly, the choice of $t_{k}$ does not depend on random information after iteration $t_{k}$.
@@
where the last inequality uses $\phi(\cdot)<0.4$. From the choice of $t_{k}$, (84) leads to
$r_{t}\leq$ $c_{\tau}(\beta)\frac{P_{t_{k}}({\bm{x}}_{t_{k}+1})}{P_{t_{k}}({\bm{x}}^{*})}\left[\frac{2M_{f}}{k}+\frac{\beta^{1/2}_{t}}{k}\sqrt{C_{\gamma}t\gamma_{t}^{f}}+(0.4+\beta^{1/2})\sigma_{t_{k}}^{f}({\bm{x}}_{t_{k}+1})\right],$ (85)
with probability $\geq 1-\delta$. Next, we consider the probability function $P_{t_{k}}$ at $\bm{x}$$^{*}$ and ${\bm{x}}_{t_{k}+1}$. Using the fact that $c({\bm{x}}^{*})\leq 0$, we have by Lemma C.2 at $\bm{x}$$^{*}$ and $t_{k}$,
Thus, we can write
$\frac{-\mu_{t_{k}}^{c}({\bm{x}}^{*})}{\sigma_{t_{k}}^{c}({\bm{x}}^{*})}\geq-B_{c}.$ (87)
From the monotonicity of $\Phi$, we have
$\Phi\left(\frac{-\mu_{t_{k}}^{c}({\bm{x}}^{*})}{\sigma_{t_{k}}^{c}({\bm{x}}^{*})}\right)\geq\Phi\left(-B_{c}\right),$ (88)
@@
$\frac{P_{t_{k}}({\bm{x}}_{t_{k}+1})}{P_{t_{k}}({\bm{x}}^{*})}\leq\frac{1}{\Phi(-B_{c})}.$ (89)
Applying (89) to (85), we have
$r_{t}\leq$ $c_{\tau}(\beta)\frac{1}{\Phi(-B_{c})}\left[\frac{2M_{f}}{k}+\frac{\beta^{1/2}_{t}}{k}\sqrt{C_{\gamma}t\gamma_{t}^{f}}+(0.4+\beta^{1/2})\sigma_{t_{k}}^{f}({\bm{x}}_{t_{k}+1})\right],$ (90)
with probability $\geq 1-\delta$.
@@
From Lemma A.5, $\sigma_{i}^{f}({\bm{x}}_{i+1})\geq\frac{\sqrt{\gamma_{t}^{f}t}}{k}$, where $i=0,\dots,t-1$, at most $k$ times for any $k\in\mathbb{N}$ and $k\leq t$. Choose $k=[t/3]$ which leads to $3k\leq t\leq 3(k+1)$. Let $t_{k}$ be the first index in $[k,3k]$ so that $\sigma_{t_{k}}^{f}({\bm{x}}_{t_{k}+1})\leq\frac{\sqrt{t\gamma^{f}_{t}}}{k}$ and $\max\{f_{t_{k}}^{+}-\mu_{t_{k}}^{f}({\bm{x}}_{t_{k}+1}),0\}<\frac{2M_{f}}{k}+\frac{\beta^{1/2}_{t}}{k}\sqrt{C_{\gamma}t\gamma_{t}^{f}}$, which exists with probablity $\geq 1-2\delta/3$. Notice that $\beta_{t}^{1/2}=\mathcal{O}(\log^{1/2}(t))$. Following the proof for (90), we have
$r_{t}\leq$ $c_{\tau}(\beta)\frac{1}{\Phi(-B_{c})}\left[\frac{2M_{f}}{k}+\frac{\beta^{1/2}_{t}}{k}\sqrt{C_{\gamma}t\gamma_{t}^{f}}+(0.4+\beta^{1/2})\frac{\sqrt{t\gamma^{f}_{t}}}{k}\right],$ (91)
@@
Appendix D Test problems
The mathematical formulations of the testing problems in Section 4.2 are given in this section. The objective, constraint functions and the optimal $f$ of Problem 1 is given below.